Factor each polynomial by grouping the first two terms and the last two terms.
x^3-4x^2+2x-8
Thank you.

Answer by venugopalramana(3286)

x^3-4x^2+2x-8
=X^2(X-4)+2(X-4)=(X-4)(X^2+2)